### v1

# CTPU-PTC-24-02

CTPU-PTC-24-02
EPHOU-24-001
Finite modular axion and radiative moduli stabilization
[…]
${}^{a}$ Department of Physics, Keio University, Yokohama, 223-8522, Japan
${}^{b}$ Center for Theoretical Physics of the Universe, Institute for Basic Science (IBS), Daejeon 34051, Korea
${}^{c}$ Department of Physics, Hokkaido University, Sapporo 060-0810, Japan
We propose a simple setup which can stabilize a modulus field of the finite modular symmetry by the Coleman-Weinberg potential. Our scenario leads to a large hierarchy suppressing instanton-like corrections $e^{2\pi i\tau}$ and to a light axion identified as $\mathrm{Re}\tau$, where $\tau$ is the modulus field. This stabilization mechanism provides the axion solution to the strong CP problem. The potential has a minimum at a large $\mathrm{Im}\tau$ which suppresses explicit $U(1)_{\mathrm{PQ}}$ violation terms proportional to $e^{-2\pi{\mathrm{Im}\tau}}$, and hence the quality of the axion is ensured by the residual symmetry associated with the $T$-transformation, $\tau\to\tau+1$, around the fixed point $\tau\sim i\infty$.
1 Introduction
The finite modular symmetry has been intensively studied to explain the flavor structure of the Standard Model (SM) quarks and leptons [1, 2, 3, 4, 5, 6, 7, 8, 9]. (See for reviews [10, 11].) Under the symmetry, a Yukawa coupling constant behaves as the so-called modular form which is a holomorphic function of the modulus $\tau$. Interestingly, the finite modular symmetries $\Gamma_{N}$, $N\in\mathbb{N}$, are isomorphic to the non-Abelian discrete symmetries [12], which can explain the flavor structure of the SM fermions [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Such modular dependence of the Yukawa coupling to a modulus was found in heterotic orbifold models [23, 24, 25, 26, 27, 28] and magnetized D-brane models [29, 30, 31, 32, 33, 34, 35] in superstring theory. Moreover, the finite modular symmetry can explain the hierarchies of the quark and lepton masses and mixing if the modulus is stabilized near one of the fixed points where the residual symmetry remains unbroken [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47].
In the models with finite modular flavor symmetries, the value of modulus $\tau$ plays an essential role to fit to the data and is often treated as a free parameter. Whereas in a UV theory, its value should be fixed dynamically through moduli stabilization. There are studies on the moduli stabilization with the three-form fluxes [48, 49], the negative power of modular form [50, 51] 111 The positive power of the trivial singlet is also studied in Ref. [51]. See also Ref. [52]. , and the general superpotential invariant under the modular symmetry $SL(2,\mathbb{Z})$ [53, 54, 55, 56, 57], utilizing the Klein $j$ function [58]. The latter two mechanisms will need non-perturbative dynamics to generate those potentials. Recently, effect of the radiative correction to the tree-level stabilization is discussed in Ref. [59].
In this work, we point out that the modulus can be stabilized only by the Coleman-Weinberg (CW) potential generated by couplings with matter fields. This mechanism provides a perturbative way to stabilize a modulus through the supersymmetry breaking. For illustration, we shall consider a simple model with vector-like quarks which transforms as non-trivial singlets under the finite modular flavor symmetry. The resultant potential has a global minimum at $\mathrm{Im}\,\tau\gg 1$, where the residual $\mathbb{Z}^{T}_{N}$ symmetry associated with the $T$ transformation, $\tau\to\tau+1$, is unbroken. That can lead to a large hierarchy suppressing instanton-like correction, i.e., $e^{-2\pi{\mathrm{Im}\,\tau}}\ll 1$. An interesting feature of this mechanism is that the $\mathrm{Re}\,\tau$ direction can be identified as the QCD axion due to the accidental Peccei-Quinn (PQ) symmetry originated from the $\mathbb{Z}^{T}_{N}$ residual symmetry. We shall discuss the condition that the axion solution to the strong CP problem is not spoiled due to the CW potential, i.e. the axion quality is good enough.
This paper is organized as follows. The finite modular symmetry is briefly reviewed in Sec. 2. In Sec. 3, we discuss the mechanism of the modular stabilization by utilizing the CW potential, and then we argue existence of the QCD axion in the setup in Sec. 4. Section 5 concludes.
2 Finite modular symmetry
We consider the series of groups
$\Gamma(N):=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}(2,\mathbb{Z}),\quad\begin{pmatrix}a&b\\ c&d\end{pmatrix}\equiv\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\quad\mathrm{mod}~{}N\right\},$ (2.7)
where $N\in\mathbb{N}$ is called level and $\Gamma:=SL(2,\mathbb{Z})=\Gamma(1)$ is a group of $2\times 2$ matrices with determinant unity. The modular transformation $\gamma\in\Gamma(N)$ for a complex parameter $\tau$ is defined as
$\tau\to\gamma\tau=\frac{a\tau+b}{c\tau+d}.$ (2.8)
The modular group is generated by three generators
$S=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},\quad T=\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\quad R=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}.$ (2.15)
The finite modular symmetry is defined as the quotient group $\Gamma_{N}:=\overline{\Gamma}/\Gamma(N)$ where $\overline{\Gamma}:=\Gamma(1)/\mathbb{Z}^{R}_{2}$. Here $\mathbb{Z}^{R}_{2}$ is the $\mathbb{Z}_{2}$ symmetry generated by $R$. Under $\Gamma_{N}$ with $N<6$, the generators satisfy
$S^{2}=(ST)^{3}=T^{N}=1.$ (2.16)
There are Abelian discrete symmetries $\mathbb{Z}^{S}_{2}$, $\mathbb{Z}^{ST}_{3}$ and $\mathbb{Z}^{T}_{N}$ in associated with the $S$, $ST$ and $T$ generators which are unbroken at $\tau=i$, $w:=e^{2\pi i/3}$ and $i\infty$, respectively. The finite modular symmetries are isomorphic to the non-Abelian discrete symmetries, e.g. $\Gamma_{3}\simeq A_{4}$. A modular form $Y^{({k})}_{r}(\tau)$ with a modular weight $k$ and representation $r$ transforms under $\Gamma_{N}$ as
$Y^{({k})}_{r}(\tau)\to(c\tau+d)^{k}\rho(r)Y^{({k})}_{r}(\tau),$ (2.17)
where $\rho(r)$ is the representation matrix of $\Gamma_{N}$. We assume that the chiral superfield $Q$ with weight $-k_{Q}$ and representation $r_{Q}$ transforms as
$Q\to(c\tau+d)^{-k_{Q}}\rho(r_{Q})Q.$ (2.18)
More detailed discussions can be found in e.g. Refs. [10, 11].
3 Radiative moduli stabilization
We consider a simple supersymmetric model with the following Kähler potneital and superpotnetial:
$K=$ $\ -h\log(-i\tau+i\tau^{\dagger})+\sum_{i}\left(\frac{Q_{i}^{\dagger}Q_{i}}{(-i\tau+i\tau^{\dagger})^{k_{Q_{i}}}}+\frac{\overline{Q}_{i}^{\dagger}\overline{Q}_{i}}{(-i\tau+i\tau^{\dagger})^{k_{\overline{Q}_{i}}}}\right),$ (3.1)
$W=$ $\ \sum_{i}M_{Q_{i}}Y^{({k_{i}})}_{r_{i}}(\tau)\overline{Q}_{i}Q_{i},$
where $h\in\mathbb{N}$. The reduced Planck mass $M_{p}=2.4\times 10^{18}~{}\mathrm{GeV}$ is set to unity. The chiral superfield $Q_{i}$ ($\overline{Q}_{i}$) has a modular weight $k_{Q_{i}}$ ($k_{\overline{Q}_{i}}$) and is assumed to be a singlet under $\Gamma_{N}$ for simplicity. Here, $Q_{i}$ and $\overline{Q}_{i}$ are vector-like pairs under the SM gauge group. The weight of the modular form $k_{i}=k_{Q_{i}}+k_{\overline{Q}_{i}}-h$ and the representation $r_{i}$ are chosen so that the combination $e^{K}|W|^{2}$ in supergravity is invariant under $\Gamma_{N}$. We introduce the vector-like mass parameter $M_{Q_{i}}$ which could be replaced by vacuum expectation values of fields. Throughout this work, we assume that the SM fields are trivial singlets under the modular symmetry $\Gamma_{N}$ for simplicity and they have no effects in the following analysis.
We shall show that the modulus $\tau$ can be stabilized without introducing a superpotential for the modulus $\tau$, but with the 1-loop CW potential
$V_{\mathrm{CW}}=\frac{1}{32\pi^{2}}\sum_{i}$ $\left[\left(m_{i}^{2}+m_{Q_{i}}^{2}(\tau)\right)^{2}\left(\log\left(\frac{m_{i}^{2}+m_{Q_{i}}^{2}(\tau)}{\mu^{2}}\right)-\frac{3}{2}\right)\right.$ (3.2)
$\hskip 85.35826pt\left.-(m_{Q_{i}}^{2}(\tau))^{2}\left(\log\left(\frac{m_{Q_{i}}^{2}(\tau)}{\mu^{2}}\right)-\frac{3}{2}\right)\right],$
with
$m_{Q_{i}}^{2}(\tau):=M_{Q_{i}}^{2}\left(-i\tau+i\tau^{\dagger}\right)^{k_{i}}\left|{Y^{({k_{i}})}_{r_{i}}(\tau)}\right|^{2}.$ (3.3)
Here, we consider the $\overline{\mathrm{MS}}$ scheme with $\mu$ being the renormalization scale. The first (second) term in the parenthesis corresponds to the scalar (fermion) contribution. The supersymmetric mass $m_{Q_{i}}$ is multiplied by the factor $(-i\tau+i\tau^{\dagger})^{k_{i}/2}$ due to the canonical normalization. We introduce the soft supersymmetry breaking mass squared $m_{i}^{2}$ for the scalar component of $Q_{i}$ which is assumed to be independent of $\tau$ for simplicity. This is the case if supersymmetry is broken by a mechanism irrelevant to the modulus $\tau$. We shall first see the simplest case analytically at large $\mathrm{Im}\,\tau$, and then see the results numerically later. Throughout this paper, we assume that the mass parameters are common for $Q_{i}$’s, i.e. $M_{Q_{i}}=:M_{Q}$, $m_{i}=:m_{0}$ and $k_{i}=:k$ for simplicity. Note that the tree-level scalar potential in supergravity is irrelevant when $Q_{i}=\overline{Q}_{i}=0$.
3.1 Simplified analysis
We argue the model of radiative stabilization of the modulus $\tau$ where there is only one pair of $(\overline{Q},Q)$, so we omit the index $i$. The singlet modular form of $\Gamma_{N}$ can be expanded as
$Y^{({k})}_{1_{t}}(\tau)=q^{t/N}\sum_{n=0}^{\infty}c_{n}q^{n},\quad q:=e^{2\pi i\tau},$ (3.4)
where $r=1_{t}$ is a singlet whose charge is $0\leq t<N$ under the $\mathbb{Z}^{T}_{N}$ symmetry 222 In the usual notation [1], $11_{0}=$ 1, $1_{1}=1^{\prime}$ and $1_{2}=1^{{\prime\prime}}$ for $A_{4}$. . Assuming $\left|{q}\right|=e^{-2\pi\mathrm{Im}\,\tau}\ll 1$ and $m_{0}^{2}\ll M_{Q}^{2}$, the CW potential is given by
$V_{\mathrm{CW}}=\frac{m_{0}^{2}\tilde{M}_{Q}^{2}}{16\pi^{2}}x^{k}e^{-px}\left(\log\frac{\tilde{M}_{Q}^{2}}{e\mu^{2}}+k\log x-px\right)+\mathcal{O}\left({\left|{q}\right|,m_{0}^{4}}\right),$ (3.5)
where $x:=2\mathrm{Im}\,\tau$, $p:=2\pi t/N$ and $\tilde{M}_{Q}:=c_{0}M_{Q}$. At the leading order, there is no potential for $\mathrm{Re}\,\tau$ which will be important for the axion interpretation discussed in the next section. The first derivative of the potential is given by
$\frac{dV_{\mathrm{CW}}}{dx}\sim\frac{m_{0}^{2}\tilde{M}_{Q}^{2}}{16\pi^{2}}x^{k-1}e^{-px}\left(k-px\right)\left(\log\frac{\tilde{M}_{Q}^{2}}{\mu^{2}}+k\log x-px\right).$ (3.6)
The point $x=k/p$ is a maximum for sufficiently large $k/p$, and hence this potential can have a minimum at
$x_{0}=-\frac{k}{p}\mathcal{W}\left(-\frac{p}{k}\left(\frac{\mu}{\tilde{M}_{Q}}\right)^{2/k}\right),$ (3.7)
where the second parenthesis of Eq. (3.6) is vanishing. Here, $\mathcal{W}(z)$ is the Lambert function satisfying $\mathcal{W}(z)e^{\mathcal{W}(z)}=z$ which has two real values for $-e^{-1}<z<0$, and the larger one can be a solution where our approximation $\left|{q}\right|\ll 1$ is valid. Taking $\mu=\tilde{M}_{Q}$ and $p=2\pi/3~{}$, the location of the minimum is $x_{0}=3.9,~{}7.9,~{}12$ and $16$ for $k=6,8,10$ and $12$, respectively. $x_{0}$ increases slightly by decreasing $\mu$. Thus we find that there is the minimum at $\mathrm{Im}\,\tau\gg 1$ in the CW potential. Note that this minimum exists only for $t>0$ where the modular form $Y^{({k})}_{1_{t}}$ is a non-trivial singlet. For the CW potential from the trivial-singlet modular form, the potential increases as $\mathrm{Im}\,\tau$ increases in contrast to that from the non-trivial one, and thus there is no minimum at $\mathrm{Im}\,\tau\gg 1$ 333 The potentials discussed in Refs. [53, 54, 55, 56, 57, 58] have a similar behavior, that is, $V\to\infty$ for $\mathrm{Im}\,\tau\to\infty$. See also the potential in Ref. [52]. Therefore the dominant CW potential should be generated from the non-trivial singlet one for the existence of the minimum at $\mathrm{Im}\,\tau\gg 1$, so that the potential tends to approach to zero for a sufficiently large ${\rm Im}\tau$.
3.2 Numerical analysis
Figure 1: The shape of the CW potential along the $\mathrm{Im}\,\tau$ direction with $\mathrm{Re}\,\tau=0$ (red solid) and $\mathrm{Re}\,\tau=-0.5$ (blue dashed).
We shall study the potential Eq. (3.2) numerically without relying on the $q$-expansion in Eq. (3.4). For illustration, we consider $\Gamma_{3}\simeq A_{4}$ as the finite modular group, and two pairs of vector-like quarks whose superpotential is given by
$W=M_{Q}\sum_{i=1,2}Y^{({12})}_{1_{i}}(\tau)\overline{Q}_{i}Q_{i}.$ (3.8)
We can easily choose the representations and modular weights such that only those two terms are allowed but mixing terms like $\bar{Q}_{1}Q_{2}$ are forbidden. An explicit assignment of representations and modular weights will be shown in Eq. (4.13) when we shall discuss the vector-like quark decays. The modular forms of weight 12 are given by
$Y^{({12})}_{1_{1}}(\tau)=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}(Y_{3}^{2}+2Y_{1}Y_{2}),\quad Y^{({12})}_{1_{2}}(\tau)=(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})^{2},$ (3.9)
where the functions $Y_{1,2,3}$ are defined as [1]
$Y_{1}(\tau)=$ $\ \frac{i}{2\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}-27\frac{\eta^{\prime}(3\tau)}{\eta(3\tau)}\right],$ (3.10)
$Y_{2}(\tau)=$ $\ \frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+w^{2}\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+w\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right],$ (3.11)
$Y_{3}(\tau)=$ $\ \frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+w\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+w^{2}\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right],$ (3.12)
with $\eta(\tau)$ being the Dedekind eta function. The $q$-expansion of the modular forms are given by 444 There is an ambiguity of the normalization of the modular forms which can not be determined from the symmetry. We simply employ the normalization used in Ref. [1]. See also Ref. [60].
$Y^{({12})}_{1_{1}}=-12q^{1/3}\left(1+472q+\mathcal{O}\left({q^{2}}\right)\right),\quad Y^{({12})}_{1_{2}}=144q^{2/3}\left(1+224q+\mathcal{O}\left({q^{2}}\right)\right).$ (3.13)
The modular weights are chosen to $k=12$ so that the strong CP problem is solved with keeping the modulus mass heavy, as discussed in the next section 555 Since $1_{1}$ and $1_{2}$ can be in the $2^{\prime}$ representation under $A_{4}^{\prime}$, the two fields could be embedded into a doublet. . Here, we introduce the two vector-like pairs, so that the mixed anomaly of $A_{4}$ with the QCD is canceled [61, 62]. The singlets are trivial under the $S$-transformation, and transform as $Y^{({k})}_{1_{t}}\to w^{t}Y^{({k})}_{1_{t}}$ under the $T$-transformation. Hence, in our setup, $\det(\rho(T))=w^{1+2}=1$ and the anomaly is canceled. The anomaly associated with the modular weight can be canceled by the Green-Schwarz mechanism with an another gauge coupling modulus or by adding $\log Y^{({k})}_{1_{0}}(\tau)$ to the gauge kinetic function with a certain coefficient, where $Y^{({k})}_{1_{0}}(\tau)$ is a trivial singlet modular form and could be explained by threshold corrections from heavy modes [63].
Figure 1 shows the shape of the potential along the $\mathrm{Im}\,\tau$ direction. Here, we take $m^{2}/M_{Q}^{2}=10^{-8}$ and $\mu/M_{Q}=10^{-2}$ so that the minimum resides at $\mathrm{Im}\,\tau\simeq 13$. The red solid (blue dashed) line corresponds to $\mathrm{Re}\,\tau=0$ ($-0.5$). Note that the fundamental domain is $\mathrm{Im}\,\tau\geq 1$ ($\sqrt{3}/2$) at $\mathrm{Re}\,\tau=0~{}(-0.5)$. In the left panel, the origin of the horizontal axis is $\sqrt{3}/2$ and the global picture of the potential is shown. The right panel is the shape of the potential around the global minimum at large $\mathrm{Im}\,\tau$. For $\mathrm{Im}\,\tau\lesssim 1$, where the $q$-expansion is not efficient, there is a local minimum at $\tau\sim w$, but $V_{\mathrm{CW}}(\tau=w)=0$, since $Y^{({12})}_{1_{1,2}}(w)=0$, is shallower than the global minimum at $\mathrm{Im}\,\tau\sim 12$ where the potential value is negative. The potential clearly depends on $\mathrm{Re}\,\tau$ at small $\mathrm{Im}\,\tau$, while it is independent of $\mathrm{Re}\,\tau$ within numerical precision. Thus, the CW potential in the $\mathrm{Re}\,\tau$ direction is approximately flat near the global minimum at $\mathrm{Im}\,\tau\gg 1$. As a result, our scenario of modulus stabilization can lead a light axion as well as a large hierarchy by $|q|=e^{-2\pi\mathrm{Im}\,\tau}\sim 10^{-36}$. We will confirm this flatness analytically in the next section to discuss the quality of the axion solution to the strong CP problem.
4 Axion solution and its quality
We have seen that the CW potential Eq. (3.2) induced by the superpotential Eq. (3.8) is very flat in the $\mathrm{Re}\,\tau$ direction. The stabilization mechanism can realize the axion solution to the strong CP problem by assuming that the matter fields are vector-like quarks as in the KSVZ axion model [64, 65]. The effective $\theta$-angle $\overline{\theta}$ in the QCD is given by
$\overline{\theta}(\tau)=\theta_{0}+\mathrm{Arg}\left(Y^{({12})}_{1_{1}}(\tau)Y^{({12})}_{1_{2}}(\tau)\right)=\theta_{0}+\phi+\mathcal{O}\left({\left|{q}\right|}\right),$ (4.1)
where $\theta_{0}$ is a constant and $\phi:=2\pi\mathrm{Re}\,\tau$. The $\phi$ dependence appears as a result of the chiral anomaly of the QCD with the approximate $U(1)_{\rm PQ}$, which originates from the residual $\mathbb{Z}^{T}_{3}$ symmetry as shown below. For $\mathrm{Im}\,\tau\gg 1$, the Yukawa coupling is given by
$Y^{({k})}_{1_{t}}\overline{Q}_{t}Q_{t}\propto e^{i(t/3)\phi}\;\overline{Q}_{t}Q_{t},$ (4.2)
so we can find an accidental $U(1)_{\mathrm{PQ}}$ symmetry as
$\phi\to\phi+\alpha,\quad\overline{Q}_{t}Q_{t}\to e^{-it\alpha/3}\overline{Q}_{t}Q_{t},$ (4.3)
where $\alpha$ is a real transformation parameter. Hence the PQ charge of $\overline{Q}_{t}Q_{t}$ is $-t$, and the $U(1)_{\rm PQ}$ is anomalous. This symmetry is ensured from the residual $\mathbb{Z}^{T}_{3}$ symmetry corresponding to $\alpha=2\pi$, and the explicit breaking effects of the $U(1)_{\mathrm{PQ}}$ are from $q=e^{2\pi i\tau}$ which is invariant under $\mathbb{Z}^{T}_{3}$ but violates $U(1)_{\mathrm{PQ}}$ 666 The corrections to the axion from $\log Y^{({k})}_{1_{0}}$ in the gauge kinetic function which may exist to cancel the weight anomaly will also appear at $\mathcal{O}\left({q}\right)$. . Therefore, $\mathrm{Re}\,\tau$ can be identified as the QCD axion. Since the kinetic term of the modulus is given by 777 We write explicitly the reduced Planck mass $M_{p}$ in this section.
$\frac{hM_{p}^{2}}{(2\mathrm{Im}\,\tau)^{2}}\partial_{\mu}\tau^{\dagger}\partial^{\mu}\tau,$ (4.4)
the axion decay constant is $f_{a}=\sqrt{h}M_{p}/(2\pi x_{0})\sim\mathcal{O}\left({10^{16}}\right)$ GeV which requires a fine-tuning of the initial condition or early matter domination [66], to avoid the overabundance of the axion.
The axion potential is given by
$V_{a}=-\Lambda_{\mathrm{QCD}}^{4}\cos\left(\overline{\theta}(\phi)\right)+\Delta V.$ (4.5)
Here, $\Delta V:=V_{\mathrm{CW}}|_{x=x_{0}}$ is the axion-dependent part of the CW potential which can spoil the axion solution. The shift of the $\theta$-angle due to $\Delta V$ is estimated as
$\Delta\theta=\frac{1}{\Lambda_{\mathrm{QCD}}^{4}}\left(\frac{\partial\Delta V}{\partial\phi}\right)\simeq\frac{m_{0}^{2}M_{Q}^{2}}{8\pi^{2}\Lambda_{\mathrm{QCD}}^{4}}x_{0}^{k}\epsilon^{4+N}\Omega\sin{\theta_{0}},$ (4.6)
with
$\Omega:=\left(\beta_{1}\frac{kN-4\pi x_{0}}{kN-2\pi x_{0}}-\beta_{2}\right)\left(\log\frac{m^{2}_{0}+m_{Q_{2}}^{2}}{e\mu^{2}}+\frac{m_{Q_{2}}^{2}}{m_{0}^{2}}\log\left(1+\frac{m^{2}_{0}}{m_{Q_{2}}^{2}}\right)\right),$ (4.7)
where $\epsilon:=e^{-\pi x_{0}/N}$. In our model, $k=12$ and $N=3$. Here, $m^{2}_{Q_{2}}$ is the mass squared of $Q_{2}$ defined in Eq. (3.3) with taking $x=x_{0}$ and dropping the negligible corrections at $\mathcal{O}\left({q}\right)$. $\beta_{t}$ is the ratio $c_{1}/c_{0}$ in the expansion of the modular forms $Y^{({k})}_{1_{t}}$ in Eq. (3.4), which is given by $\beta_{1}=472$ and $\beta_{2}=224$ for $Y^{({12})}_{1_{1}}$ and $Y^{({12})}_{1_{2}}$, respectively. Note that the $\mathcal{O}\left({q}\right)$ correction proportional to $m_{Q_{1}}^{2}$ is canceled at the minimum $x=x_{0}$ and thus the leading term is $\epsilon^{4+N}$ which is proportional to $m_{Q_{2}}^{2}\sim\mathcal{O}\left({\epsilon^{4}x_{0}^{k}M_{Q}^{2}}\right)$. This shift of the angle should be less than $\mathcal{O}\left({10^{-10}}\right)$ to be consistent with the measurement of the neutron electric dipole moment [67]. A possible modulus $\tau$-dependence of $\Lambda_{\mathrm{QCD}}$ from the gauge kinetic function will also be suppressed by $\mathcal{O}\left({q}\right)$, and thus is sub-dominant for the stabilization. Similarly the $\phi$ dependence of $\partial_{\phi}\overline{\theta}$ is also negligible. In the $A_{4}$ model discussed in the previous section, we obtain
$\left|{\Delta\theta}\right|\simeq 2\times 10^{-10}\times\sin\theta_{0}\left(\frac{\left|{\Omega}\right|}{10^{4}}\right)\left(\frac{m_{0}}{10^{7}~{}\mathrm{GeV}}\right)^{2}\left(\frac{M_{Q}}{M_{p}}\right)^{2}\left(\frac{x_{0}}{28}\right)^{8}\left(\frac{\epsilon}{10^{-12}}\right)^{7}.$ (4.8)
Thus the axion quality is ensured if $\epsilon\lesssim 10^{-12}$, where $\mathrm{Im}\,\tau\sim 13$ as shown in Fig. 1. The required tuning becomes mild if $\sin\theta_{0}\ll 1$ which might happen when the (generalized) CP is conserved at some level as discussed in e.g. Ref. [56]  [^8] See Refs. [68, 69] for recent studies on the strong CP problem relevant to the modular symmetry.. In such a case, the strong CP problem is partially solved by the modular axion $\phi\sim\mathrm{Re}\,\tau$.
Figure 2: Masses of the modulus $X\sim\mathrm{Im}\,\tau$ (red) and vector-like quarks $Q_{1}$ (blue dashed) and $Q_{2}$ (green dot-dashed). In this plot, $m_{0}=10^{7}~{}\mathrm{GeV}$, $M_{Q}=10^{4}\times\mu=M_{p}$, and $h=3$.
Before closing, we discuss the masses of the modulus $X\propto\mathrm{Im}\,\tau$ and the vector-like quarks $Q_{1,2}$. The modulus mass may need to be heavier than $\mathcal{O}\left({10~{}\mathrm{TeV}}\right)$, so that the modulus decay occurs before the big-bang nucleosynthesis and the moduli problem is avoided [70, 71, 72, 73]. After canonically normalizing the kinetic term Eq. (4.4), the modulus mass is given by
$m_{X}^{2}=\frac{m_{0}^{2}\tilde{M}_{Q}^{2}}{8\pi^{2}hM_{p}^{2}}x_{0}^{k}\epsilon^{2}(k-px_{0})^{2}.$ (4.9)
Hence the modulus mass is related to $\Delta\theta$ as
$m_{X}\simeq$ $\ c_{0}\left|{k-px_{0}}\right|\sqrt{\frac{\Delta\theta}{h\Omega\sin{\theta_{0}}}}\frac{\Lambda_{\mathrm{QCD}}^{2}}{\epsilon^{1+N/2}M_{p}}$ (4.10)
$\sim$ $\ 42~{}\mathrm{TeV}\times\frac{c_{0}\left|{k-2\pi x_{0}/N}\right|}{\sqrt{h\sin{\theta_{0}}}}\left(\frac{\Delta\theta}{10^{-10}}\right)^{1/2}\left(\frac{10^{4}}{\Omega}\right)^{1/2}\left(\frac{\Lambda_{\mathrm{QCD}}}{100~{}\mathrm{MeV}}\right)^{2}\left(\frac{10^{-12}}{\epsilon}\right)^{5/2},$
in the $A_{4}$ model. Thus the modulus mass decreases as the axion quality is higher by $\epsilon\to 0$. For $\Delta\theta/\sin\theta_{0}\sim 10^{-10}$, $\epsilon\lesssim 10^{-12}$ is required to avoid the moduli problem 999 The lighter modulus would not be a problem if there is the thermal inflation [74, 75, 76], large Hubble induced mass during the inflation exists [77, 78], and/or an inflation scale is lower than the modulus mass. .
The mass of the lighter vector-like quark $Q_{2}$, whose $T$ charge is $2$, is given by
$m_{Q_{2}}\sim\tilde{M}_{Q}x_{0}^{k/2}\epsilon^{2}\sim 1~{}\mathrm{TeV}\times\left(\frac{\tilde{M}_{Q}}{M_{p}}\right)\left(\frac{x_{0}}{28}\right)^{6}\left(\frac{\epsilon}{10^{-12}}\right)^{2},$ (4.11)
thus $M_{Q}\sim M_{p}$ is necessary for $m_{Q_{2}}\gtrsim\mathcal{O}\left({\mathrm{TeV}}\right)$ to be consistent with the LHC constraints [79, 80, 81, 82, 83]. Conversely, if we require $m_{Q_{2}}\gtrsim\mathrm{TeV}$ and $\tilde{M}_{Q}\lesssim M_{p}$, the hierarchy parameter is bounded from below as $\epsilon\gtrsim 10^{-8}\times x_{0}^{-k/4}$, and thus larger weight is preferred to allow $\epsilon\sim\mathcal{O}\left({10^{-12}}\right)$ without having the too light vector-like quark. Figure 2 shows the masses of the modulus $X\propto\mathrm{Im}\,\tau$ and the vector-like quarks. As $\mathrm{Im}\,\tau$ increases, the axion quality becomes better, whereas the modulus and vector-like quark masses decrease, and hence there is the $lower$ bound $\Delta\theta\gtrsim 10^{-13}$ for sufficiently heavy vector-like quark $Q_{2}$.
The vector-like quarks should be able to decay into SM particles. For illustration, if the vector-like quarks have the same quantum number as the down-type quarks, denoted by $d$, we can write down the superpotential
$W_{\mathrm{mix}}=\sum_{i=1,2}{\lambda}_{i}Y^{({14})}_{1_{1}}H_{d}qQ_{i},$ (4.12)
where $H_{d}$ and $q$ are respectively the down-type Higgs doublet and the doublet SM quark. Here, the representations and weights of quarks are set to
$r_{q}=r_{d}=r_{\overline{Q}_{1}}=1_{0},\quad r_{Q_{1,2}}=r_{\overline{Q}_{2}}=1_{2},\quad k_{q}=h-k_{d}=8,\quad k_{Q_{1,2}}-h=k_{\overline{Q}_{1,2}}=6,$ (4.13)
and those of $H_{d}$ is the trivial-singlet with weight 00. We omit the flavor indices of the SM quarks. With this assignment, the interactions in Eqs. (3.8) and (4.12), as well as the SM Yukawa coupling $yH_{d}qd$ with $y$ being a constant, are realized, while the $\overline{Q}_{i}d$ is vanishing because of the negative weight. The size of mixing $s_{Q_{i}}$ is estimated as
$s_{Q_{i}}\sim$ $\ \frac{{\lambda}_{i}x_{0}^{7}Y^{({14})}_{1_{1}}{\langle{H_{d}}\rangle}}{M_{Q}x_{0}^{6}Y^{({12})}_{1_{i}}}\sim\frac{{\lambda}_{i}x_{0}{\langle{H_{d}}\rangle}}{M_{Q}\epsilon^{i-1}}$ (4.14)
$\sim$ $\ 10^{-3}\times{\lambda}_{2}\left(\frac{x}{28}\right)\left(\frac{{\langle{H_{d}}\rangle}}{100~{}\mathrm{GeV}}\right)\left(\frac{M_{p}}{M_{Q}}\right)\left(\frac{10^{-12}}{\epsilon}\right),$
where $i=2$ in the second line, and we used $Y^{({k})}_{1_{t}}\sim\epsilon^{t}$. Thus the vector-like quarks decay promptly at the collider scale. It is noted that the CW potential induced by the mixing term is negligible because ${\langle{H_{d}}\rangle}\ll M_{Q}$.
5 Summary and discussions
In this work, we point out that the modulus of the finite modular symmetry can be stabilized by the Coleman-Weinberg (CW) potential. For illustration, we studied the model with $\Gamma_{3}\simeq A_{4}$ symmetry and two vector-like quark pairs which transform as non-trivial singlets, namely $1_{1}$ and $1_{2}$. This is the minimal possibility to cancel the mixed QCD anomaly of the finite modular symmetry, but to induce that of the $U(1)_{\mathrm{PQ}}$ symmetry. The CW potential has the global minimum at $\mathrm{Im}\,\tau\gg 1$ if the corresponding modular form is non-trivial singlet under $A_{4}$, where the residual symmetry $\mathbb{Z}^{T}_{3}$ remains unbroken. Since the potential along the $\mathrm{Re}\,\tau$ direction has extremely flat due to this residual symmetry, we can regard $\mathrm{Re}\,\tau$ as the QCD axion to solve the strong CP problem. The accidental $U(1)_{\mathrm{PQ}}$ arises due to the $\mathbb{Z}^{T}_{3}$ symmetry and hence the PQ breaking effects are controlled by $\Gamma_{3}$ which is a different situation argued in Ref. [84]. We examined the condition to ensure the quality of this finite modular axion, and correlate with the masses of the modulus $X\sim\mathrm{Im}\,\tau$ and vector-like quarks. Interestingly, the modulus $X$ and the lighter quark $Q_{2}$ are expected to be $\mathcal{O}\left({\mathrm{TeV}}\right)$ scale for $\Delta\theta<10^{-10}$ and thus could be probed by cosmology and collider experiments. This mechanism can be generalized to other modular forms as long as the CW potential is dominated by that with non-zero $T$-charge and the potential converges to zero for $\mathrm{Im}\,\tau\gg 1$.
The modular $A_{4}$ symmetry in our scenario can not be used to explain the hierarchical structure of the SM fermions, as studied in Refs. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], since the hierarchy parameter $\epsilon$ is $\mathcal{O}\left({10^{-12}}\right)$ and is too small. Such a tiny $\epsilon$ is required due to the correlation of the axion quality $\Delta\theta$ with the modulus mass in Eq. (4.10). This relation can be relaxed if the $\tau$-independent angle $\theta_{0}$ is small or the level $N$ is large. In the latter case, there exist a small hierarchy $|e^{2\pi i\tau/N}|$ in the SM sector and a large one $|e^{2\pi i\tau}|$ in the vector-like quark sector.101010 In general, for example, if one could break the periodicity $\tau\sim\tau+1$ to $\tau\sim\tau+1/n$ ($n>1$) in the vector-like $(Q,\bar{Q})$ sector with keeping the periodicity $\tau\sim\tau+1$ in the SM fermion sector, then modular forms may appear $Y^{k}_{r}(\tau)$ in the SM fermion sector and $Y^{k^{\prime}}_{r^{\prime}}(n\tau)$ in the $(Q,\bar{Q})$ sector, and one could realize both a small hierarchy $|e^{2\pi i\tau}|$ in the SM sector and a large one $|e^{2\pi in\tau}|$ in the vector-like quark pairs. . Thus, we could construct a model which explains the SM fermion hierarchies by the modulus stabilized by the mechanism proposed in this paper. An explicit model building is subject of our future work.
We have proposed a new scenario to lead to a large hierarchy $e^{-2\pi\mathrm{Im}\,\tau}\sim\mathcal{O}\left({10^{-36}}\right)$ and a light axion by stabilizing the modulus at $\mathrm{Im}\,\tau\gg 1$. Although we have applied it to solve the strong CP problem by assuming $Q$ and $\bar{Q}$ as vector-like matter fields with the QCD colors, our scenario would be useful to explain other large hierarchies, which would be required for the proton stability, tiny neutrino masses, CP/flavor violations, quintessence and so on, together with light axions by assuming $Q$ and $\bar{Q}$ as visible or hidden matter fields. We would study it elsewhere.
In this work, we assume that the soft scalar masses of the vector-like quarks are independent of the modulus $\tau$ and supersymmetry is dominantly broken by other sector, so that the supersymmetry breaking sector does not contribute to the stabilization of the modulus $\tau$. In addition, there would be contributions from gravitational instantons [85] which would change the potential structure. Those effects and other dynamical effects including stabilization mechanisms of other moduli, and uplifting of the vacuum energy if necessary, may become important in complete models such as superstring, but beyond the scope of this paper.
Acknowledgement
We thank Y.Abe and N.Aso for helpful discussions in the early stage of this work. The work of J.K. is supported in part by the Institute for Basic Science (IBS-R018-D1). This work is supported in part by he Grant-in-Aid for Scientific Research from the Ministry of Education, Science, Sports and Culture (MEXT), Japan No. JP22K03601 (T.H.) and JP23K03375 (T.K.).